# ProFL: Performative Robust Optimal Federated Learning

### v1

@@ 5 Performance Analysis
To facilitate analysis, we consider scenarios where $D_{i}(\theta)=\sum_{k=1}^{K}\nu_{i,k}(\theta)D_{i,k}(\theta)$ represents a mixture of $K$ group distributions. $\nu_{i,k}(\theta)$ is the proportion of samples from group $k$ and $\sum_{k=1}^{K}\nu_{i,k}(\theta)=1$. Let $f_{i,k}(.)$ denote the function for each group’s distribution.
[…]
@@
Assumption 5.4 (PL Condition Polyak, (1964)).
@@ 5.1 Error of the performative gradient
Under Assumption 5.3, let $\lambda_{i,\min}$ denote the minimal singular value of $\Delta\theta_{i}$ defined in Eq. (6) during all iterations. With probability at least $1-\varphi_{i}$, the estimation error of $\frac{\widehat{df_{i}}}{d\theta}$ is bounded by $\boldsymbol{\omega_{F}}$ specified as follows:
@@
Under Assumptions 5.1, 5.2 and 5.3, the following holds with probability at least $1-\varphi_{i}$:
@@
Additionally, the notations $\boldsymbol{\omega_{D}}$ and $\boldsymbol{\omega_{Q}}$ represent upper bounds. From Lemmas 5.8 and 5.9, we find that since the estimation error of performative gradient is primarily dominated by the estimation error of $\nabla\mathcal{L}_{i,2}$, we have:
@@ 5.2 Convergence analysis
Given Lemmas 5.7, 5.8, and 5.9, we are now ready to analyze the convergence. Denote $\overline{\gamma}=\sum_{i\in\mathcal{V}}\alpha_{i}\gamma_{i}$, $\sigma^{2}_{\gamma}=\sum_{i\in\mathcal{V}}\alpha_{i}(\overline{\gamma}-\gamma_{i})^{2}$, $\overline{\epsilon}=\sum_{i\in\mathcal{V}}\alpha_{i}\epsilon_{i}$, and $\varphi=\sum_{i\in\mathcal{V}}\varphi_{i}.$ Throughout the training process, let $W_{1}(D,Q)_{\max}$ represent the maximum Wasserstein-1 distance $W_{1}\left(D(\theta_{i}^{t}),Q_{i}\right)$ between actual data distribution and contamination $\forall i\in\mathcal{V}$. Let $\overline{\boldsymbol{\omega_{Q}}},\overline{\boldsymbol{\omega_{D}}}$ be the maximum of $\boldsymbol{\omega_{Q}}$ and $\boldsymbol{\omega_{D}}$ respectively in Lemma 5.9 $\forall i\in\mathcal{V}$. Let $g^{2}_{\min}$ be the minimal value of $\mathbb{E}[\|g^{t}\|^{2}]$ with $g^{t}=\sum_{i\in\mathcal{V}}\alpha_{i}\sum_{k=1}^{K}\widehat{\nabla}\mathcal{L}_{{i,k}}.$
@@
Let $\{\overline{\theta}^{t}\}_{t\geq 0}$ be a sequence of global models generated by ProFL (and PoFL). Under Assumptions 5.1, 5.2, 5.3, and 5.4, with probability at least $1-T\varphi$, we have:
$\mathbb{E}\left[\mathcal{L}\left(\overline{\theta}^{T}\right)-\mathcal{L}\left({\theta}^{\text{PO}}\right)\right]\leq$ $\left(1-\rho\eta\right)^{T}\mathbb{E}\left[\mathcal{L}\left(\overline{\theta}^{0}\right)-\mathcal{L}\left({\theta}^{\text{PO}}\right)\right]+\left((1-\overline{\epsilon})\overline{\boldsymbol{\omega_{D}}}^{2}+\overline{\epsilon}\left(\overline{\boldsymbol{\omega_{D,Q}}}\right)^{2}-\frac{g^{2}_{\min}}{2}\right)\eta$
$\quad+\left(L(1+\overline{\gamma})\left((1-\overline{\epsilon})\overline{\boldsymbol{\omega_{D}}}^{2}+\overline{\epsilon}\left(\overline{\boldsymbol{\omega_{D,Q}}}\right)^{2}+G^{2}\right)\right)\eta^{2}$
$\quad+\frac{1}{2}L^{2}(R-1)^{2}G^{2}\left((1+\overline{\gamma})^{2}+\sigma^{2}_{\gamma}\right)\eta^{3}.$
@@
Theorem 5.10, together with Lemmas 5.7, 5.8 and 5.9 in Section 5.1, highlight the impact of sampling error, data contamination, client heterogeneity, and the non-linearity of $f_{i}(\theta)$ on convergence, as discussed below.
@@
The term $L^{2}(R-1)^{2}G^{2}\eta\left((1+\overline{\gamma})^{2}+\sigma^{2}_{\gamma}\right)$ captures the effect of client heterogeneity. Recall that $R$ is the number of local updates. In the special case where $R=1$ , meaning clients update and aggregate models at each iteration, this term becomes zero.
@@
$\mathbb{E}\left[\mathcal{L}(\overline{\theta}^{T})-\mathcal{L}({\theta}^{\text{PO}})\right]\leq$ $(1-\rho\eta)^{T}\mathbb{E}\left[\mathcal{L}(\overline{\theta}^{0})-\mathcal{L}({\theta}^{\text{PO}})\right]+\left(\overline{\boldsymbol{\omega_{D}}}^{2}-\frac{g^{2}_{\min}}{2}\right)\eta+L(1+\overline{\gamma})\left(\overline{\boldsymbol{\omega_{D}}}^{2}+G^{2}\right)\eta^{2}$
$\quad+\left(\frac{1}{2}L^{2}(R-1)^{2}G^{2}\left((1+\overline{\gamma})^{2}+\sigma^{2}_{\gamma}\right)\right)\eta^{3}.$
@@
  1. 1.
    If $g^{2}_{\min}>0$, we can always adjust learning rate $\eta$, estimation window $H$, number of local updates $R$, and sample size $n_{i}$ such that $\overline{\boldsymbol{\omega_{D}}}^{2}<g^{2}_{\min}/2$ and the following holds
@@
  2. 2.
    If $g^{2}_{\min}=0$, this means there is at least one iteration $t$ that has $\|g^{t}\|=0$ and $\mathbb{E}[\mathcal{L}(\overline{\theta}^{t+1})-\mathcal{L}({\theta}^{\text{PO}})]=\mathbb{E}[\mathcal{L}(\overline{\theta}^{t})-\mathcal{L}({\theta}^{\text{PO}})]$. Except for the iterations where $\|g^{t}\|=0$, all the other iterations have positive $\|g^{t}\|>0$. We can choose $\eta,H,n_{i}$ such that $\overline{\boldsymbol{\omega_{D}}}^{2}<g^{2}_{\min}/2$ and the following holds
@@
    where $b$ is the number of iterations with $\|g^{t}\|=0$.
All above theoretical results apply to both ProFL and PoFL. The robust strategies in ProFL reduce the upper bound in Theorem 5.10.
@@ 5.3 Discussion
Section 3 introduced simple yet effective strategies in ProFL to enhance PoFL. We now explain why these strategies are effective.
[…]
@@ 6 Experiments
Experimental Setup. We evaluate our algorithm in five case studies: 1) pricing with dynamic demands, 2) pricing with dynamic contribution, 3) binary strategic classification, 4) house pricing regression, and 5) regression with dynamic contribution.
Pricing with Dynamic Demands. Consider a company that aims to find the proper prices for various goods. Let $\theta\in\mathbb{R}^{d}$ be the prices associated with the set of $d$ goods, and $Z_{i}\in\mathbb{R}^{d}$ be the retailer $i$’s demands for these goods. We assume $Z_{i}\sim\mathcal{N}(\mu(\theta),\sigma^{2})$ is model dependent Gaussian distribution, i.e., mean demand for each good decreases linearly as the price increases. $\mu(\theta)=\mu_{0}-\gamma_{i}\theta$, here $\gamma_{i}\geq 0$ measures the price sensitivity and varies across different retailers. The goal is to maximizes total revenue over all retailers $\sum_{i=1}^{N}\alpha_{i}\mathbb{E}_{Z_{i}\sim D_{i}(\theta)}[\theta^{T}Z_{i}]$.
Pricing with Dynamic Contribution. Consider a similar setting as above where a company aims to find prices for various goods that maximize the total revenue. Suppose the retailers it faces have consumers from $K$ groups (e.g., old/young) and these consumers have fixed demands but they have options to purchase the goods from other companies. Here we assume the demands of retailer $i$’s consumers from group $k\in[K]$ have fixed distribution $D_{i,k}=\mathcal{N}(\mu_{i,k},\sigma_{i}^{2})$, but the fraction of groups change dynamically based on prices $\theta$. Let $r_{i,k}(\theta)=B-\mathbb{E}_{Z_{i}\sim D_{i,k}}[\theta^{T}Z_{i}]$ be the expected remaining budget of group $k$ after purchasing the goods (with initial budget $B$) and assume the fraction of group $k$ is $\nu_{i,k}(\theta)=\frac{r_{i,k}(\theta)}{\sum_{k=1}^{K}r_{i,k}(\theta)}$ (i.e., groups with more remaining budgets have more incentives to stay in the system). Then the retailer $i$’s total demands would follow the mixture of distributions $D_{i}(\theta)=\sum_{k=1}^{K}\nu_{i,k}(\theta)D_{i,k}$.
Binary Strategic Classification. Consider an FL system is used to make binary decisions about human agents (e.g., loan application, hiring, college admission), where client $i$ with data $Z_{i}=(X_{i},Y_{i})$ may manipulate their features strategically to increasing their chances of receiving favorable outcomes at the lowest costs. Suppose the clients are assigned positive decisions whenever $21/\left(1+\exp(-X^{T}\theta)\right)\geq 1/$ 2 and the cost it takes to manipulate feature from $X_{i}$ to $X_{i}^{\prime}$ is $\frac{1}{2\gamma_{i}}||X^{\prime}-X||$, then strategic clients would react to the decision model $\theta$ by manipulating their features to $\arg\max_{X_{i}^{\prime}}\left(-\langle X^{\prime}_{i},\theta\rangle-\frac{1}{2\gamma_{i}}\|X_{i}^{\prime}-X_{i}\|_{2}^{2}\right)=X_{i}-\gamma_{i}\theta$. For the initial data, we acquire it from a synthetic data where $X|Y=y\sim\mathcal{N}(\mu_{y},\sigma_{y}^{2})$ for $y\in\{0,1\}$ and two realistic Adult and Give Me Some Credit data. The details of these data are in Appendix E.1. The goal is to minimizes prediction loss over all clients.
House Pricing Regression. Consider a regression task that aims to find the proper listing prices for houses based on features such as size, age, number of bedrooms. Suppose the features of houses in district $i$ follow $X_{i}\sim\mathcal{N}(\mu_{i},\sigma_{i}^{2})$. Given predictive model $\theta$, let the listing price be $h(X_{i})=\theta^{T}X_{i}$. Since the houses with higher listing prices tend to lower the demand, the actual selling price $Y$ depends on $\theta$ and we assume $Y_{i}=(a_{i}-\gamma_{i}\theta)^{T}X_{i}+\epsilon$, where $\epsilon\sim\mathcal{N}(0,\sigma^{2}_{n})$ is Gaussian noise. The goal is to minimizes prediction error
Regression with dynamic contribution. Consider an example of retail inventory management, where the regression model $\theta$ predicts product sales considering features like seasonal trends and advertising. Distributor $i$ has retailers from two groups $k\in\{1,2\}$, and each retailer’s data distribution is fixed. The fraction of retailers from each group is $\nu_{i,k}(\theta)=\frac{\ell_{i,-k}(\theta)+c}{\sum_{k^{\prime}\in\{1,2\}}\left(\ell_{i,k^{\prime}}(\theta)+c\right)}$ where $-k=\{1,2\}\setminus k$ because the retailers are more likely to stop purchasing if they experience higher prediction error $\ell_{i,k}(\theta)$. The goal is to find $\theta$ that minimizes expected prediction error $\sum_{i=1}^{N}\alpha_{i}\mathbb{E}_{(X_{i},Y_{i})\sim D_{i}(\theta)}\left[(\theta^{T}X_{i}-Y_{i})^{2}\right]$, where $Y_{i}$ is a weighted combination of fixed distributions.
For binary classification, we use both synthetic and real data (Adult and Give Me Some Credit datasets), while the other cases use synthetic data. All dynamics are synthetic. The parameter choices is in Table 3 of Appendix E.1. Information on the real datasets is in Appendix E.4. Each method is run 10 times under the same setup, and averages and variance are reported.
@@
Baselines. To date, only Jin et al., (2024) has addressed model-dependent distribution shifts in FL. Thus, we compare our algorithm with this Performative Federated Learning (PFL). Unlike our algorithm, theirs converges to a performative stable solution ${\theta}^{\text{PS}}$ under stricter assumptions. We also compare the Performative Gradient (PG) algorithm from Izzo et al., (2021), in the centralized setting.
Table 1: Performative Loss of PG and ProFL

### v2

@@ 4 Performance Analysis
To facilitate analysis, we study two special cases: (i) Contribution dynamics, where only $\nu_{i,k}(\theta)$ changes while the group distribution remains fixed, i.e., $D_{i}(\theta)=\sum_{k\in[K]}\nu_{i,k}(\theta)D_{i,k}$. In this case, $f_{i,k}(\theta)=\nu_{i,k}(\theta)$, and $\widehat{f}_{i,k}$ estimates the sample proportion from group $k$. (ii) Distribution dynamics, where only the distribution $D_{i,k}(\theta)$ changes while the contribution from each group remains fixed, i.e., $D_{i}(\theta)=\sum_{k\in[K]}\nu_{i,k}D_{i,k}(\theta)$. We consider a mixture of Gaussians with $D_{i,k}(\theta)=\mathcal{N}(f_{k}(\theta),\sigma_{k}^{2})$, where $f_{i,k}(\theta)$ is the mean of group $k$ for client $i$, and the overall mean for client $i$ is $f_{i}(\theta)=\sum_{k\in[K]}\nu_{i,k}f_{i,k}(\theta)$. $\widehat{f}_{i}(\mathcal{S}_{i})$ represents the empirical mean from data samples $\mathcal{S}_{i}$. Our theoretical results apply to both cases. Before presenting them, we introduce the technical assumptions. For simplicity, $\|.\|_{2}$ is denoted as $\|.\|$. Proofs are in Appendix.
@@
Assumption 4.4 (PL Condition polyak1964some).
@@ 4.1 Error of the performative gradient
Under Assumption 4.3, let $\lambda_{i,\min}$ denote the minimal singular value of $\Delta\theta_{i}$ defined in Eq. (3.2.1) during all iterations. With probability at least $1-\varphi_{i}$, the estimation error of $\frac{\widehat{df_{i}}}{d\theta}$ is bounded by $\boldsymbol{\omega_{F}}$ specified as follows:
@@
Under Assumptions 4.1, 4.2 and 4.3, the following holds with probability at least $1-\varphi_{i}$:
@@
Additionally, the notations $\boldsymbol{\omega_{D}}$ and $\boldsymbol{\omega_{Q}}$ represent upper bounds. From Lemmas 4.8 and 4.9, we find that since the estimation error of performative gradient is primarily dominated by the estimation error of $\nabla\mathcal{L}_{i,2}$, we have:
@@ 4.2 Convergence analysis
Given Lemmas 4.7, 4.8, and 4.9, we are now ready to analyze the convergence. Denote $\overline{\gamma}=\sum_{i\in\mathcal{V}}\alpha_{i}\gamma_{i}$, $\sigma^{2}_{\gamma}=\sum_{i\in\mathcal{V}}\alpha_{i}(\overline{\gamma}-\gamma_{i})^{2}$, $\overline{\epsilon}=\sum_{i\in\mathcal{V}}\alpha_{i}\epsilon_{i}$, and $\varphi=\sum_{i\in\mathcal{V}}\varphi_{i}.$ Throughout the training process, let $W_{1}(D,Q)_{\max}$ represent the maximum Wasserstein-1 distance $W_{1}\left(D(\theta_{i}^{t}),Q_{i}\right)$ between actual data distribution and contamination $\forall i\in\mathcal{V}$. Let $\overline{\boldsymbol{\omega_{Q}}},\overline{\boldsymbol{\omega_{D}}}$ be the maximum of $\boldsymbol{\omega_{Q}}$ and $\boldsymbol{\omega_{D}}$ respectively in Lemma 4.9 $\forall i\in\mathcal{V}$. Let $g^{2}_{\min}$ be the minimal value of $\mathbb{E}[\|g^{t}\|^{2}]$ with $g^{t}=\sum_{i\in\mathcal{V}}\alpha_{i}\sum_{k=1}^{K}\widehat{\nabla}\mathcal{L}_{{i,k}}.$
@@
Let $\{\overline{\theta}^{t}\}_{t\geq 0}$ be a sequence of global models generated by ProFL (and PoFL). Under Assumptions 4.1, 4.2, 4.3, and 4.4, with probability at least $1-T\varphi$, we have:
$\mathbb{E}\left[\mathcal{L}\left(\overline{\theta}^{T}\right)-\mathcal{L}\left({\theta}^{\text{PO}}\right)\right]\leq$
$~\left(1-\rho\eta\right)^{T}\mathbb{E}\left[\mathcal{L}\left(\overline{\theta}^{0}\right)-\mathcal{L}\left({\theta}^{\text{PO}}\right)\right]$
$+\left((1-\overline{\epsilon})\overline{\boldsymbol{\omega_{D}}}^{2}+\overline{\epsilon}\left(\overline{\boldsymbol{\omega_{D,Q}}}\right)^{2}-\frac{g^{2}_{\min}}{2}\right)\eta$
$+\left(L(1+\overline{\gamma})\left((1-\overline{\epsilon})\overline{\boldsymbol{\omega_{D}}}^{2}+\overline{\epsilon}\left(\overline{\boldsymbol{\omega_{D,Q}}}\right)^{2}+G^{2}\right)\right)\eta^{2}$
$+\frac{1}{2}L^{2}(R-1)^{2}G^{2}\left((1+\overline{\gamma})^{2}+\sigma^{2}_{\gamma}\right)\eta^{3}.$
@@
Theorem 4.10, together with Lemmas 4.7, 4.8 and 4.9 in Section 4.1, highlight the impact of sampling error, data contamination, client heterogeneity, and the non-linearity of $f_{i}(\theta)$ on convergence, as discussed below.
@@
The term $L^{2}(R-1)^{2}G^{2}\eta\left((1+\overline{\gamma})^{2}+\sigma^{2}_{\gamma}\right)$ captures client heterogeneity. Recall that $R$ is the number of local updates. In the special case where $R=1$ , meaning clients update and aggregate models at each iteration, this term becomes zero.
@@
$\mathbb{E}\left[\mathcal{L}(\overline{\theta}^{T})-\mathcal{L}({\theta}^{\text{PO}})\right]\leq(1-\rho\eta)^{T}\mathbb{E}\left[\mathcal{L}(\overline{\theta}^{0})-\mathcal{L}({\theta}^{\text{PO}})\right]$
$+\left(\overline{\boldsymbol{\omega_{D}}}^{2}-\frac{g^{2}_{\min}}{2}\right)\eta+L(1+\overline{\gamma})\left(\overline{\boldsymbol{\omega_{D}}}^{2}+G^{2}\right)\eta^{2}$
$+\left(\frac{1}{2}L^{2}(R-1)^{2}G^{2}\left((1+\overline{\gamma})^{2}+\sigma^{2}_{\gamma}\right)\right)\eta^{3}.$
@@
1) If $g^{2}_{\min}>0$, we can always adjust learning rate $\eta$, estimation window $H$, number of local updates $R$, and sample size $n_{i}$ such that $\overline{\boldsymbol{\omega_{D}}}^{2}<g^{2}_{\min}/2$ and the following holds
@@
2) If $g^{2}_{\min}=0$, this means there is at least one iteration $t$ that has $\|g^{t}\|=0$ and $\mathbb{E}[\mathcal{L}(\overline{\theta}^{t+1})-\mathcal{L}({\theta}^{\text{PO}})]=\mathbb{E}[\mathcal{L}(\overline{\theta}^{t})-\mathcal{L}({\theta}^{\text{PO}})]$. Except for the iterations where $\|g^{t}\|=0$, all the other iterations have positive $\|g^{t}\|>0$. We can choose $\eta,H,n_{i}$ such that $\overline{\boldsymbol{\omega_{D}}}^{2}<g^{2}_{\min}/2$ and the following holds
@@
where $b$ is the number of iterations with $\|g^{t}\|=0$. All above theoretical results apply to both ProFL and PoFL. The robust strategies in ProFL reduce the upper bound in Theorem 4.10.
@@ 4.3 Discussion
Handling complex distribution shifts with non-linear $f_{i}(\theta)$. Based on Lemma 4.7, the non-linear function $f_{i}(\theta)$ primarily impacts the estimation error through the term $\frac{M^{2}\eta^{2}G^{4}H^{6}}{2}$, which decreases as $\eta$ andHH are reduced. However, to ensure the matrix $\Delta\theta_{i}$ in Eq. (3.2.1) is nonsingular, H1H-1 must be larger than the dimensionality of $\theta$. In cases where $H$ cannot be decreased, reducing $\eta$ is essential to attain a smaller $\boldsymbol{\omega_{F}}$. Moreover, the upper bound in Theorem 4.10 takes the form $A\eta+B\eta^{2}+C\eta^{3}$, where $B,C,\eta>0$. Thus, a smaller $\eta$ results in a smaller upper bound. However, reducing $\eta$ will also increase the error term $\frac{2f_{\varphi}(\varphi_{i};n_{i})Hd}{n_{i}\eta^{2}}$ in Eq. (8). Nevertheless, if the sample size $n_{i}$ is sufficiently large and $\frac{M^{2}\eta^{2}G^{4}H^{6}}{2}\gg\frac{2f_{\varphi}(\varphi_{i};n_{i})Hd}{n_{i}\eta^{2}}$, reducing the learning rate is still effective, as we verify in Fig. 3(b). In cases where local samples are limited, estimating $\widehat{\frac{df_{i}}{d\theta}}$ at the server side could be considered, e.g., by clustering clients with similar performative distribution map $D_{i}$ and perform global aggregation to mitigate estimation error.
@@ 5 Experiments
5.1 Experimental Setup.
Case Studies. We evaluate our algorithm on five case studies covering both types of dynamics introduced in Section 4: (i) Contribution dynamics, where group contributions change dynamically; and (ii) Distributional dynamics, where group distributions change dynamically. We adopt the same mixture-model formulation $D_{i}(\theta)$ as in the theoretical analysis.
Specifically, we study five scenarios covering both dynamics: (1) Dynamic Pricing with Demand Shift—prices influence mean demands across retailers; (2) Dynamic Pricing with Group Contributions—group shares of demand adjust based on remaining budgets; (3) Binary Strategic Classification—clients strategically manipulate features to maximize favorable outcomes; (4) House Pricing Regression—listing and selling prices interact with the model’s predictions; and (5) Regression with Dynamic Contributions—group fractions of retailers adapt to model prediction errors. Scenarios (2) and (5) correspond to Contribution dynamics, while the remaining scenarios correspond to Distributional dynamics.
Client Configuration and Heterogeneity. For each experimental setting, every method is independently run 10 times with random seeds 0–9. We report the average and variance across these runs. Table 1 and Figures 23 (case study: pricing with dynamic demands) use 10 heterogeneous clients, except the $\alpha=0$ case in Table 1, which uses homogeneous clients, and Figure 2(b), which uses 100 clients. Figures 4 (binary strategic classification), 6 (house pricing regression), 5(a) (pricing with dynamic contribution), and 5(b) (regression with dynamic contribution) all use 10 heterogeneous clients.
@@
Baselines. To date, only jin2023performative has addressed model-dependent distribution shifts in FL. Thus, we compare our algorithm with this Performative Federated Learning (PFL). Unlike our algorithm, theirs converges to a performative stable solution ${\theta}^{\text{PS}}$ under stricter assumptions. We also compare with the Performative Gradient (PG) algorithm from pmlr-v139-izzo21a in the centralized setting, as well as a straightforward extension of PG to the FL setting without our robust strategies, which we refer to as PoFL (Performative Optimal Federated Learning).
Dataset. Binary classification uses both synthetic and real data (Adultmisc_adult_2 and Give Me Some CreditGiveMeSomeCredit datasets), while the other cases use synthetic data. All dynamics are synthetic.
The supplementary material provides complete descriptions of the case study settings, data-generation models, training configurations, and all parameter values used in the experiments.
5.2 Numerical Results
Comparison to Centralized Setting. Table 1 shows that adapting PG to the FL framework improves performance. $\gamma_{i}\in[\overline{\gamma}-\alpha,\overline{\gamma}+\alpha]$. Larger $\alpha$ indicates more heterogeneity. As client heterogeneity increases, the performance of both algorithms degrades. Under homogeneous distribution shifts ($\alpha=0$), both algorithms perform similarly. ProFL outperforms PG at $\alpha=\{0.25,0.5\}$, since the centralized method estimates only one $\frac{df}{d\theta}$ on the server and fails under variance shifts.